A box of books weighing 304 N moves with
a constant velocity across the floor when it is
pushed with a force of 373 N exerted downward at an angle of 39.4◦ below the horizontal.
Find µk between the box and the floor.
1
Expert's answer
2019-12-17T09:54:24-0500

The FBD is shown below



The Newton's second law in projections on the axes


"F_f=F\cos\theta""N=mg+F\sin\theta"

The friction force


"F_f=\mu_kN"

Hence


"\mu_k=\frac{F\cos\theta}{mg+F\sin\theta}""\mu_k=\frac{373\cos 39.4^{\circ}}{304+373\sin 39.4^{\circ}}=0.533"